20 December 2007

RHIC Startup:  Recovered from Maintenance Day with first store at 9:00PM.  Began with a 6x6 bunch store, with an 87x87 bunch store for physics at 10:40PM. 

Using the new dAu81 lattice with beta* of 0.8 in Yellow.  The new lattice appears to have a smaller emittance in Yellow than Blue at the beginning of the store.

STAR lost their BBC counters and needed to make an access (circa 5:00AM) to swap a card, so they do not have an estimate of the background rates with the new lattice.  Vacuum near STAR is improved somewhat, down to 1E-10, down from 2E-10.  PHENIX reports that the backgrounds look somewhat better.  Travis tuning the Blue collimators had to pull them out from their set-points to reduce the PHENIX backgrounds reported by scintillator S3.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020